# Piezoelectric actuation for integrated photonics

### v1

@@ Abstract
Recent decades have seen significant advancements in integrated photonics, driven by improvements in nanofabrication technology. This field has developed from integrated semiconductor lasers and low-loss waveguides to optical modulators, enabling the creation of sophisticated optical systems on a chip scale capable of performing complex functions like optical sensing, signal processing, and metrology. The tight confinement of optical modes in photonic waveguides further enhances the optical nonlinearity, leading to a variety of nonlinear optical phenomena such as optical frequency combs, second-harmonic generation, and supercontinuum generation. Active tuning of photonic circuits is crucial not only for offsetting variations caused by fabrication in large-scale integration, but also serves as a fundamental component in programmable photonic circuits. Piezoelectric actuation in photonic devices offers a low-power, high-speed solution and is essential in the design of future photonic circuits due to its compatibility with materials like Si and Si3N4, which do not exhibit electro-optic effects. Here, we provide a detailed review of the latest developments in piezoelectric tuning and modulation, by examining various piezoelectric materials, actuator designs tailored to specific applications, and the capabilities and limitations of current technologies. Additionally, we explore the extensive applications enabled by piezoelectric actuators, including tunable lasers, frequency combs, quantum transducers, and optical isolators. These innovative ways of managing photon propagation and frequency on-chip are expected to be highly sought after in the future advancements of advanced photonic chips for both classical and quantum optical information processing and computing.
@@ 1 Introduction
Integrated photonics [1, 2] has helped bring table-top optical systems in the research lab to the chip scale with compact sizes, portable weight, and low power consumption. Thanks to the advancement of nanofabrication technologies, significant achievements have been made in integrated semiconductor lasers [3, 4, 5, 6, 7, 8], modulators [9, 10, 11, 12, 13, 14, 15, 16], photodetectors [17, 18], as well as ultralow-loss PICs [19, 20, 21, 22, 23, 24, 25, 26]. Moreover, the high confinement of light in integrated optical waveguides allows the exploration of the optical nonlinearities [27, 28], giving rise to nonlinear photonics such as optical frequency comb (OFC) generation [29], supercontinua [30], and second harmonic generation [31, 32]. For instance, dissipative Kerr soliton (DKS) microcombs formed in Kerr-nonlinear optical microring resonators [29, 28, 33] has provided a miniaturized, coherent, broadband OFC source with repetition rates in the millimeter-wave to microwave domain, which has found wide applications from atomic clock [34, 35], telecommunication [36], spectroscopy [37], to Light Detection and Ranging (LiDAR) [38], as well as the heterogeneous integration with III-V/Si lasers [39, 40, 41, 7].
@@
Efficient and high-speed tuning of integrated photonic devices is becoming inevitable for many advanced applications. For example, fast actuation of microring resonators allows stabilizing the microcomb repetition rate with high bandwidth [34, 54]. Also, resonance tuning can be used for tunable filters and compensating fabrication errors. Moreover, the programmability of photonic circuits enables more complex photonic circuits to be built, such as photonic neural network [55] and photonic quantum computing [56]. On the other hand, microwave frequency (Gigahertz, GHz) optical modulation is highly demanded recently in platforms such as spatio-temporal modulation based optical non-reciprocity [57] and topological optical band structures in synthetic frequency dimensions [58, 59, 60], which could open up new avenues for studying novel physics phenomenon.
@@
Conventionally, the optical waveguides can be modulated by incorporating electro-optic materials (e.g., LiNbO3, AlN), whose refractive index can be changed upon external electric field, originating from their nonlinearity. However, for materials such as Si and Si3N4, due to their inversion symmetry, and thus the lack of nonlinearity, it is difficult to electrically modulate them. Over a long time, the thermo-optic tuning is utilized [61, 62, 63, 64, 37, 55], which is limited by slow tuning speed (1 ms), high power consumption (1 mW), and large thermal cross-talk. Therefore, it is incompatible with large-scale integration and cryogenic applications [65] due to these disadvantages. Although electro-optic modulation of Si waveguides through the free-carrier injection [66, 67] and carrier depletion effect [68, 69] has made promising progress, it suffers from large propagation loss due to free carrier absorption. Alternatively, heterogeneous or hybrid integration with electro-optic materials, e.g., graphene [70, 71, 72, 73], monolayer WS2 [74], lead zirconate titanate (PZT) [75], lithium niobate (LiNbO3) [10, 76, 77, 78, 11, 79, 80], barium titanate (BaTiO3) [81, 82, 83, 84], and organic electro-optic materials [85], has been intensively studied recently. However, there are still open questions with regard to CMOS-compatibility, fabrication complexity, optical losses, and dispersion engineering.
@@
The stress-optic effect has gained significant attention for the actuation of passive photonic waveguides and microring resonators due to the advances in Micro-Electro-Mechanical Systems (MEMS) technologies [86, 87]. By applying external forces either piezoelectrically [88, 89, 12, 90, 91, 92] or electrostatically [93, 94, 95, 96, 97], the deformation of the mechanical structure will generate stress and strain on the optical waveguides, and thus change the effective refractive index of the waveguide through photoelastic [98] and moving boundary effects [99]. Although electrostatic actuation has been successfully applied to mechanically bend photonic waveguides and resonators in various applications, e.g. directional coupler [100], optical switches [101, 102, 103], and phase shifters [104], there remain challenges that need to be addressed in the future. For example, the optical waveguides have to be suspended, which may compromise their robustness, and the air-cladded waveguides present higher optical losses. In addition, the tuning is quadratic with applied voltage, and actuation bandwidth is limited by the fundamental mechanical resonance on the order of MHz [105]. On the contrary, the piezoelectric method circumvents these problems by fully cladding the waveguide in dielectric materials. Efficient, linear, bi-directional piezoelectric tuning of Si [106] and Si3N4 [88, 89, 12, 107] microring resonators has been experimentally demonstrated, with sub-s tuning speed and nW on-hold power consumption. Moreover, acoustic waves can be excited piezoelectrically at several GHz, which enables chip-scale acousto-optic modulation. Therefore, stress-optic tuning can complement the traditional electro-optic and thermo-optic tuning in terms of material universality, CMOS compatibility, and power efficiency.
@@
This review focuses on the piezoelectric control of integrated photonic devices by giving comprehensive discussions on the aspects of piezoelectric materials, mechanical structures, and the operation frequency range. In Section 2, the fundamental physics will be introduced from piezoelectric effect to optomechanical interactions. Different mechanical modes and material platforms will be discussed. Section 3 reviews the state-of-the-art piezoelectric control of photonic integrated circuits with either released or unreleased mechanical structures, and compares the tuning performance of different piezoelectric materials. The applications of piezoelectric tuning will be elaborated in Section 4, from integrated semiconductor lasers to tunable optical frequency combs. In Section 5, acousto-optic modulation (AOM) at microwave frequency will be introduced regarding to different mechanical modes, with the applications of AOM discussed in Section 6. Section 7 will give our perspectives on current opportunities and challenges for the development of piezoelectric-controlled integrated photonic devices in the future. Finally, in Section 8, we will summarize the main features and conclude with implications for future applications. Figure 1 gives an outlook of an integrated photonic chip that is equipped with various piezoelectric actuators for different functionalities. The design of these devices will be discussed separately in each of the corresponding chapters.
@@ 2.1 Piezoelectric Effect
Piezoelectricity is an effect where charges can be generated and accumulated at the surface of the material under external pressure. It can happen only in crystalline materials with no inversion symmetry, such as PZT, AlN, and LiNbO3. The inverse effect can also happen where stress and strain are built as external electric field is applied across the material. The direct and inverse piezoelectric effect can be mathematically described by relating the electric field to the strain of the material:
where and are the strain and stress, and are the electric displacement and electric field, is the compliance coefficient that describes the microscopic Hooke’s law in linear elastic material, is the permittivity of the dielectric, is the piezoelectric strain coefficient in the unit of (which is also equal to ). Note that the piezoelectric stress coefficient (in the unit of ) is also often quoted in the literature, which characterizes the surface charge density generated by the strain. and are generally related through the stiffness coefficient of the material.
The subscript indices represent different components of each physical quantity. The capital indices are for stress and strain ranging from 1 to 6 to take into account the normal and shear components with the Voight notation that: . The indices are for electric field and displacement, and each has three components: . Note that due to the symmetry, . The piezoelectric coefficient is related to crystal structures of each piezoelectric material. Although has 18 components, due to the crystal symmetry, only few are non-zero.
@@
The piezoelectric coefficients of typical piezoelectric materials are summarized in Table 1, together with their mechanical properties. Note the specific values may vary, depending on the composition and growth conditions of the material. However, it is a good starting point for numerical simulations. For most demonstrated piezoelectric tuning of photonic circuits to date, PZT and AlN are the two main piezoelectric materials that are widely employed, because of their high e33 and maturity of foundry processes. Although PZT shows a piezoelectric coefficient nearly 20 times larger, AlN advances in better linearity, smaller dielectric constant (thus smaller capacitance and shorter charge-discharge time), and zero hysteresis (since PZT has ferroelectricity). Also, due to the high loss tangent at the microwave frequency, the actuation frequency of PZT is limited to below 1 GHz. On the other hand, AlN is excellent at microwave frequency, and efficient acousto-optic modulation (AOM) has been demonstrated [132, 14, 133, 134, 135]. Therefore, depending on the specific requirement of applications (high efficiency, linearity, or speed), the right piezoelectric material should be considered via careful design and simulation.
Scandium-doped AlN (AlScN) has attracted much attention due to its much improved piezoelectric coefficient e33 compared to AlN. Our recent attempt on the application of AlScN will be presented later. As for the LiNbO3, while itself is widely adopted as an electro-optic material, its application in piezoelectric tuning is less explored. However, its application at microwave frequencies by exciting acoustic waves for the AOM has been extensively studied due to its much higher electromechanical conversion efficiency [15, 136, 137, 138, 139, 140]. For the other piezoelectric materials Zinc Oxide (ZnO) and Lithium Tantalate (LiTaO3), which show similar properties as AlN and LiNbO3, they are still new in the field of piezoelectric tuning and AOM of photonic circuits, and further researches in this direction could potentially open up new possibilities from the perspective of integration, efficiency, and cost.
@@ 2.2 Photoelastic and Moving Boundary Effect
The generated stress and strain will modify the effective refractive index of the optical waveguide through the stress-optic effect, which consists mainly of photoelastic and moving boundary effects as will be elaborated in the following. These effects describe the interaction between light and mechanical structures (acoustic waves), which are widely adopted in fields including cavity optomechanics [141], stimulated Brillouin scattering [142], and acousto-optic modulation (AOM) [14, 15]. They offer operation in a wide frequency range from direct current (DC) to radio-frequency (RF). While the photoelastic and moving boundary effects describe the action of mechanics on the light, the inverse effects where the photon generates optical forces on the mechanical structures will also take place, which correspond to the electrostrictive [143] and radiation-pressure forces [144], respectively.
@@
where the numerator is the perturbation of the electromagnetic energy in the optical waveguide, and the denominator is the total energy of the optical mode. is the effective refractive index of the waveguide. The volume integral is over the entire region where the electromagnetic field resides. is a rank-2 tensor and the integrand in the numerator can be expanded as [15]:
where is the vacuum permittivity. The can be directly calculated from the strain-optic relation in Eq. 4. From Eqs. 10-11, we can estimate the change of the refractive index of the optical waveguide by simulating the optical and strain field distribution using Finite-Element Method (FEM).
@@
LiNbO3 | 2.21 (o) 2.14 (e) | 1.810-19 | p11=-0.026 p12=0.09 p13=0.133 p14=-0.075 p31=0.179 p33=0.071 p41=-0.151 p44=0.146 | r13=9.6 r22=6.8 r33=30.9 r51=32.6 | [98, 110, 151] |
LiTaO3 | 2.119 (o) 2.123 (e) | 1.4610-19 | p11=-0.081 p12=0.081 p13=0.093 p14=-0.026 p31=0.089 p33=-0.044 p41=-0.085 p44=0.028 | r13=8.4 r22=-0.2 r33=30.5 r51=20 | [98, 110, 151] |
BaTiO3 | 2.29 (o) 2.26 (e) | p11=0.5 p12=0.106 p13=0.2 p31=0.07 p33=0.77 p44=1.0 | r13=8 r33=40.6 r51=730 | [121, 152] | |
GaAs | 3.38 | 2.610-17 | p11=-0.165 p12=-0.140 p44=-0.072 | r41=1.43 | [110, 153, 151] |
GaP | 3.05 | 1.210-17 | p11=-0.23 p12=-0.13 p44=-0.10 | r41=-1.1 | [98, 125, 154, 155, 151] |
@@
where and are the original optical resonant angular frequency and the effective refractive index of the waveguide. The second term on the right-hand side is the relative change of the radius of the optical microring resonator due to the mechanical deformation, which can be replaced by the relative change of the cavity length for the race-track resonator. Note this is the change of the optical path which is along the longitudinal direction, and it should not be confused with the moving boundary effect which is at the transverse direction.
@@
The above equations will be useful for simulating the tuning of the optical resonance upon quasi-DC voltages applied to the actuator. For AOM at microwave frequencies, it is widely adopted to describe the interaction between photon and phonon quantum mechanically by using the single photon-phonon optomechanical coupling rate , which can be calculated by normalizing Eq. 13 using the Zero-Point Fluctuation (ZPF) of the mechanical motion as:
@@
where , and are the zero-point motion, effective mass, and resonant frequency of the mechanical resonator, respectively. is the material’s density, and and are the maximum displacement and its square over the whole mechanical mode. Intuitively, measures the scattering rate between a single photon and phonon in an optomechanical cavity, and is related with the overlap between the optical and mechanical modes. It can be experimentally calibrated by optomechanically measuring the thermal Brownian noise of the mechanical resonator [159]. From the expressions, we can also see the larger the mechanical mass and frequency , the smaller the , and thus smaller coupling rate. In the case where large is desired, it is preferred to tightly confine the mechanical mode in a small volume.
@@
The photoelastic coefficients and the optical properties of typical optical materials for integrated photonics are summarized in Table 2. The first three rows are materials with inversion symmetry, which have neither piezoelectric nor electro-optic effects. However, they are the most widely used optical materials for optical waveguides, showing ultra-low optical losses, high third-order nonlinearity, and large-scale production. On the other hand, they exhibit reasonable amount of photoelastic coefficient, which is one of the main reasons for piezoelectric tuning. For the piezoelectric materials presented in Table 1, they are also good electro-optic materials and have been applied in electro-optic modulators [160, 10]. On the other hand, they can be modulated by acoustic waves excited through their piezoelectricity by fabricating interdigital transducers (IDT) on the same layer (see next section). Most recently, efficient AOMs have been demonstrated on these material platforms [14, 15]. GaAs and GaP have recently attracted much attention in piezo-optomechanics by working as the optomechanical crystals due to their high refractive index and photoelastic coefficients [154, 161, 128, 162, 153, 163]. Sub-wavelength confinement of both optical and mechanical modes largely enhances the optomechanical coupling rate g0. Meanwhile, the high third-order nonlinearity makes them promising for nonlinear photonics. Therefore, these different materials show their own advantages in one or more aspects, and a hybrid combination is usually preferred to take full advantage.

### v2

@@ Abstract
Recent decades have seen significant advancements in integrated photonics, driven by improvements in nanofabrication technology. This field has developed from integrated semiconductor lasers and low-loss waveguides to optical modulators, enabling the creation of sophisticated optical systems on a chip-scale capable of performing complex functions like optical sensing, signal processing, and metrology. The tight confinement of optical modes in photonic waveguides further enhances the optical nonlinearity, leading to a variety of nonlinear optical phenomena such as optical frequency combs, second-harmonic generation, and supercontinuum generation. Active tuning of photonic circuits is crucial not only for offsetting variations caused by fabrication in large-scale integration, but also serves as a fundamental component in programmable photonic circuits. Piezoelectric actuation in photonic devices offers a low-power, high-speed solution and is essential in the design of future photonic circuits due to its compatibility with materials like Si and Si3N4, which do not exhibit electro-optic effects. Here, we provide a detailed review of the latest developments in piezoelectric tuning and modulation, by examining various piezoelectric materials, actuator designs tailored to specific applications, and the capabilities and limitations of current technologies. Additionally, we explore the extensive applications enabled by piezoelectric actuators, including tunable lasers, frequency combs, quantum transducers, and optical isolators. These innovative ways of managing photon propagation and frequency on-chip are expected to be highly sought after in the future advancements of advanced photonic chips for both classical and quantum optical information processing and computing.
@@ 1 Introduction
Integrated photonics [1, 2] has helped bring table-top optical systems in the research lab to the chip-scale with compact sizes, portable weight, and low power consumption. Thanks to the advancement of nanofabrication technologies, significant achievements have been made in integrated semiconductor lasers [3, 4, 5, 6, 7, 8], modulators [9, 10, 11, 12, 13, 14, 15, 16], photodetectors [17, 18], as well as ultralow-loss photonic integrated circuits (PICs) [19, 20, 21, 22, 23, 24, 25, 26]. Moreover, the high confinement of light in integrated optical waveguides allows the exploration of the optical nonlinearities [27, 28], giving rise to nonlinear photonics such as optical frequency comb (OFC) generation [29], supercontinua [30], and second harmonic generation [31, 32]. For instance, dissipative Kerr soliton (DKS) microcombs formed in Kerr-nonlinear optical microring resonators [29, 28, 33] has provided a miniaturized, coherent, broadband OFC source with repetition rates in the millimeter-wave to microwave domain, which has shown promise in widespread application domains ranging from atomic clocks [34, 35], telecommunication [36], spectroscopy [37], to Light Detection and Ranging (LiDAR) [38], as well as the heterogeneous integration with III-V/Si lasers [39, 40, 41, 7].
@@
Efficient and high-speed tuning of integrated photonic devices is becoming inevitable for many advanced applications. For example, fast actuation of microring resonators allows stabilizing the microcomb repetition rate with high bandwidth [34, 54]. Also, resonance tuning can be used for tunable filters, compensating fabrication variation, and aligning optical resonances between optical resonators [55, 56]. Moreover, the programmability of photonic circuits enables more complex photonic circuits to be built, such as photonic neural network [57] and photonic quantum computing [58]. On the other hand, microwave frequency optical modulation is highly demanded recently in platforms such as spatio-temporal modulation based optical non-reciprocity [59] and topological optical band structures in synthetic frequency dimensions [60, 61, 62], which could open up new avenues for studying novel physics phenomenon.
@@
Conventionally, the optical waveguides can be modulated by incorporating electro-optic materials, such as Lithium Niobate (LiNbO3) and Aluminum Nitride (AlN), whose refractive index can be changed upon external electric field, originating from their nonlinearity. However, for materials such as Si and Si3N4, due to their inversion symmetry, and thus the lack of nonlinearity, it is difficult to electrically modulate them. Hence, the thermo-optic tuning is utilized [63, 64, 65, 66, 37, 57], which is limited by slow tuning speed (1 ms), high power consumption (1-100 mW), and large thermal cross-talk. Therefore, it is incompatible with both large-scale photonic integration and cryogenic applications [67] that are prevalent for classical and quantum photonic computation. Although electro-optic modulation of Si waveguides through the free-carrier injection [68, 69] and carrier depletion effect [70, 71] has made promising progress, it suffers from large propagation loss due to free carrier absorption. Alternatively, heterogeneous or hybrid integration with electro-optic materials, e.g., graphene [72, 73, 74, 75], monolayer WS2 [76], lead zirconate titanate (PZT) [77], LiNbO3 [10, 78, 79, 80, 11, 81, 82], barium titanate (BaTiO3) [83, 84, 85, 86], and organic electro-optic materials [87], has been intensively studied recently. However, there are still open questions with regard to optical losses and dispersion engineering as well as the complexity of CMOS-complexity of these materials.
@@
The stress-optic effect has gained significant attention for the actuation of passive photonic waveguides and microring resonators due to the advances in Micro-Electro-Mechanical Systems (MEMS) technologies [88, 89]. By applying external forces either piezoelectrically [90, 91, 12, 92, 93, 94] or electrostatically [95, 96, 97, 98, 99], the deformation of the mechanical structure will generate stress and strain on the optical waveguides, and thus change the effective refractive index of the waveguide through photoelastic [100] and moving boundary effects [101]. Although electrostatic actuation has been successfully applied to mechanical bending of photonic waveguides and resonators in various applications, e.g. directional coupler [102], optical switches [103, 104, 105], and phase shifters [106], there remain challenges that need to be addressed in the future. For example, the optical waveguides have to be suspended, which may compromise their robustness, and the air-cladded waveguides present higher optical losses. In addition, the tuning is quadratic with applied voltage, and actuation bandwidth is limited by the fundamental mechanical resonance on the order of kHz-MHz [107]. On the contrary, the piezoelectric method circumvents these problems by fully cladding the waveguide in dielectric materials. Efficient, linear, bi-directional piezoelectric tuning of Si [108] and Si3N4 [90, 91, 12, 109] microring resonators has been experimentally demonstrated, with sub-s tuning speed and nW on-hold power consumption. Moreover, acoustic waves can be excited piezoelectrically at several GHz, which enables chip-scale acousto-optic modulation. Therefore, stress-optic tuning can complement the traditional electro-optic and thermo-optic tuning in terms of material universality, CMOS compatibility, and power efficiency.
@@
This review focuses on the piezoelectric control of integrated photonic devices by giving comprehensive discussions on the aspects of piezoelectric materials, mechanical structures, and the operation frequency range. In Section 2, the fundamental physics will be introduced from piezoelectric effect to optomechanical interactions. Different mechanical modes and material platforms will be discussed. Section 3 reviews the state-of-the-art piezoelectric control of photonic integrated circuits with either released or unreleased mechanical structures, and compares the tuning performance of different piezoelectric materials. The applications of piezoelectric tuning will be elaborated in Section 4, from integrated semiconductor lasers to tunable optical frequency combs. In Section 5, acousto-optic modulation (AOM) at Radio-Frequency (RF) will be introduced regarding to different mechanical modes, with the applications of AOM discussed in Section 6. Section 7 will give our perspectives on current opportunities and challenges for the development of piezoelectric-controlled integrated photonic devices in the future. Finally, in Section 8, we will summarize the main features and conclude with implications for future applications. Figure 1 gives an outlook of an integrated photonic chip that is equipped with various piezoelectric actuators for different functionalities. The design of these devices will be discussed separately in each of the corresponding chapters.
@@ 2.1 Piezoelectric Effect
Piezoelectricity is an effect where charges can be generated and accumulated at the surface of the material under external pressure. It can happen only in crystalline materials with no inversion symmetry, such as PZT, AlN, and LiNbO3. The inverse effect can also happen where stress and strain are built as external electric field is applied across the material. The direct and inverse piezoelectric effect can be mathematically described by relating the electric field to the strain of the material [110]:
where and are the strain and stress, and are the electric displacement and electric field, is the elastic compliance tensor that describes the microscopic Hooke’s law in linear elastic material, is the permittivity of the dielectric, is the piezoelectric strain coefficient in the unit of (which is also equal to ). Note that the piezoelectric stress coefficient (in the unit of ) is also often quoted in the literature, which characterizes the surface charge density generated by the strain. and are generally related through the elastic stiffness tensor of the material through [110]:
Subscript indices represent different components of each physical quantity. The capital indices are for stress and strain ranging from 1 to 6 to take into account the normal and shear components with the Voight notation that: . The indices are for electric field and displacement, and each has three components: . Note that due to the symmetry, . The piezoelectric coefficient is related to crystal structures of each piezoelectric material. Although has 18 components, due to the crystal symmetry, only few are non-zero.
@@
The piezoelectric coefficients of typical piezoelectric materials are summarized in Table 1, together with their mechanical properties. Note the specific values may vary, depending on the composition and growth conditions of the material. However, it is a good starting point for numerical simulations. For most demonstrated piezoelectric tuning of photonic circuits to date, PZT and AlN are the two main piezoelectric materials that are widely employed, because of their high e33 and maturity of foundry processes. Although PZT shows a piezoelectric coefficient nearly 20 times larger, AlN advances in better linearity, smaller dielectric constant (thus smaller capacitance and shorter charge-discharge time), and zero hysteresis (since PZT is ferroelectric). Also, due to the high loss tangent at microwave frequencies, the actuation frequency of PZT is limited to below 1 GHz. On the other hand, AlN is excellent at microwave frequencies, and efficient gigahertz AOM has been demonstrated [141, 14, 142, 143, 144]. Therefore, depending on the specific requirement of applications (high efficiency, linearity, or speed), the right piezoelectric material should be considered via careful design and simulation.
Scandium-doped AlN (AlScN) has attracted much attention due to its much improved piezoelectric coefficient e33 compared to AlN. As for the LiNbO3, while itself is widely adopted as an electro-optic material, its application in piezoelectric tuning is less explored. However, its application at microwave frequencies by exciting acoustic waves for the AOM has been extensively studied due to its much higher electromechanical conversion efficiency [15, 145, 146, 147, 148, 149]. For the other piezoelectric materials Zinc Oxide (ZnO) and Lithium Tantalate (LiTaO3), which show similar properties as AlN and LiNbO3, they are still new in the field of piezoelectric tuning and AOM of photonic circuits, and further research in this direction could potentially open up new possibilities from the perspective of integration, efficiency, and cost.
@@ 2.2 Photoelastic and Moving Boundary Effect
The generated stress and strain will modify the effective refractive index of the optical waveguide through the stress-optic effect, which consists mainly of photoelastic and moving boundary effects as will be elaborated in the following. These effects describe the interaction between light and mechanical structures (acoustic waves), which are widely adopted in fields including cavity optomechanics [150], stimulated Brillouin scattering [151], and AOM [14, 15]. They offer operation in a wide frequency range from direct current (DC) to radio-frequency (RF). While the photoelastic and moving boundary effects describe the action of mechanics on the light, the inverse effects where the photon generates optical forces on the mechanical structures will also take place, which correspond to the electrostrictive [152] and radiation-pressure forces [153], respectively.
@@
where the numerator is the perturbation of the electromagnetic energy in the optical waveguide, and the denominator is the total energy of the optical mode. is the effective refractive index of the waveguide. The volume integral is taken over the entire region where the electromagnetic field resides. is a rank-2 tensor and the integrand in the numerator can be expanded as [15]:
where is the vacuum permittivity. The can be directly calculated from the strain-optic relation in Eq. 5. From Eqs. 11-12, we can estimate the change of the refractive index of the optical waveguide by simulating the optical and strain field distribution using Finite-Element Method (FEM).
@@
LiNbO3 | 2.21 (o) 2.14 (e) | 1.810-19 | p11=-0.026 p12=0.09 p13=0.133 p14=-0.075 p31=0.179 p33=0.071 p41=-0.151 p44=0.146 | r13=9.6 r22=6.8 r33=30.9 r51=32.6 | [100, 113, 161, 117] |
LiTaO3 | 2.119 (o) 2.123 (e) | 1.4610-19 (800 nm) | p11=-0.081 p12=0.081 p13=0.093 p14=-0.026 p31=0.089 p33=-0.044 p41=-0.085 p44=0.028 | r13=8.4 r22=-0.2 r33=30.5 r51=20 | [100, 113, 161] |
BaTiO3 | 2.297 (o) 2.268 (e) | - | p11=0.5 p12=0.106 p13=0.2 p31=0.07 p33=0.77 p44=1.0 | r13=8 r33=40.6 r51=730 | [110, 162, 163] |
GaAs | 3.38 | 2.610-17 | p11=-0.165 p12=-0.140 p44=-0.072 | r41=1.43 | [113, 164, 161, 117] |
GaP | 3.05 | 1.110-17 | p11=-0.23 p12=-0.13 p44=-0.10 | r41=-1.1 | [100, 134, 165, 166, 161] |
@@
where and are the original optical resonant angular frequency and the effective refractive index of the waveguide. The second term on the right-hand side is the relative change of the radius of the optical microring resonator due to the mechanical deformation, which can be replaced by the relative change of the cavity length for the race-track resonator. Note this is the change of the optical path along the longitudinal direction which should not be confused with the moving boundary effect in the transverse direction.
@@
The above equations will be useful for simulating the tuning of the optical resonance upon quasi-DC voltages applied to the actuator. For AOM at microwave frequencies, it is widely adopted to describe the interaction between photon and phonon by using the single photon-phonon optomechanical coupling rate , which can be calculated by normalizing Eq. 14 using the Zero-Point Fluctuation (ZPF) of the quantum ground state of the mechanical motion as:
@@
where , and are the amplitude of zero-point motion, and the effective mass and resonant frequency of the mechanical resonator, respectively. is the material’s mass density, and and are the maximum displacement and its square over the whole mechanical mode. Intuitively, measures the scattering rate between a single photon and phonon in an optomechanical cavity, and is related with the overlap between the optical and mechanical modes. It can be experimentally calibrated by optical readout of the thermal Brownian noise of the mechanical resonator [170], typically in a homodyne interferometer. From the expressions, we can also see the larger the mechanical mass and frequency , the smaller the , and thus smaller coupling rate. In the case where large is desired, it is preferred to tightly confine the optical and mechanical modes in a small volume.
@@
The photoelastic coefficients and the optical properties of typical optical materials for integrated photonics are summarized in Table 2. The first three rows are materials with inversion symmetry, which have neither piezoelectric nor electro-optic effects. However, they are the most widely used optical materials for optical waveguides, showing ultra-low optical losses, high third-order nonlinearity, and large-scale production. On the other hand, they exhibit reasonable amount of photoelastic coefficient, which is one of the main reasons for piezoelectric tuning. For the piezoelectric materials presented in Table 1, they are also good electro-optic materials and have been applied in electro-optic modulators [171, 10]. This ‘coincidence’ is because both piezoelectric and electro-optic effects have the same microscopic origin, which is due to the non-centrosymmetry of the crystal structure of these materials, such that they support spontaneous polarization that is dependent on external stress and electric field. On the other hand, they can be modulated by acoustic waves excited through their piezoelectricity by fabricating interdigital transducers (IDT) on the same layer (see next section). Most recently, efficient AOMs have been demonstrated on these material platforms [14, 15]. GaAs and GaP have recently attracted much attention in piezo-optomechanics by working as the optomechanical crystals due to their high refractive index and photoelastic coefficients [165, 172, 137, 173, 164, 174]. Sub-wavelength confinement of both optical and mechanical modes strongly enhances the optomechanical coupling rate g0. Meanwhile, the high third-order nonlinearity makes them promising for nonlinear photonics, having recently crossed the elusive threshold for continous-travelling wave Kerr parametric gain [cite new arxiv]. Therefore, these different materials show their own advantages in one or more aspects, and a hybrid combination is usually preferred to take full advantage.